The perimeter of a rectangle is 108 inches. The length exceeds the width by 8 inches. Find the length and the width.
Again please help me.

Answer by oscargut(2103)

Let L=length and W=width
then Perimeter=P=2(L+W)=108
L=W+8
so
2(2W+8)=108
4W=96
W=24
L=32